OpenStudy (anonymous):

integral sign from -1 to pi f(x) dx = 2/3 . find dx of integrals sign from -pi to 1 f(x)

OpenStudy (anonymous):

would that be -2/3?

OpenStudy (anonymous):

So you're given \[\large\int_{-1}^\pi f(x)~dx=\frac{2}{3}\] and you're asked to find \[\large\int_{-\pi}^1f(x)~dx~~?\] I'd say we don't have enough information to conclude anything.